OpenStudy (anonymous):

what is the solution of the system 5x + 6y = 20 , 9x + 11y = 32

OpenStudy (zzr0ck3r):

ok 5x+6y=20 9x+11y=32 9(5x+6y) = 9*20 -5(9x+11y)=-5*32 45x+54y=180 -45x-55y=-160 add the equations -y=20 y=-20 so 5x+6(-20)=20 5x-120=20 5x=140 x=28

OpenStudy (zzr0ck3r):

ask questions if you dont understand

OpenStudy (anonymous):

i think i get this one thanks

OpenStudy (raden):

alternative :  asks questions too, if you dont understand this :)